# supplementary

### v2

@@ 5.1 Optimal Scaling of the Coldest Temperature Level
Roberts et al. (2022) demonstrates that, at least under stylised conditions, if we can ensure that jumping between modes is possible at the coldest temperature, then the entire algorithm (suitably time-scaled) converges to a process (Walsh Brownian Motion) which itself has a mixing time which is independent of $d$. (While the exact result (Theorem 4) in that paper assumes that jumping between modes at the cold temperature always occurs, it is easy to see from the proof that it extends routinely to just requiring that jumps happen with probability which are bounded away from 00.) Therefore this section will be dedicated to demonstrating that at a sufficiently cold inverse temperature (we need $\beta=\mathcal{O}(d)$) then the probability of accepting a move to a different mode in the cold temperature is bounded away from 00.
@@ 5.1.1 Assumptions on the Target Density
where $\sigma_{j}\in\mathbb{R}_{+}$, $\mu_{ij}\in\mathbb{R}$ and $f$ is an un-normalised univariate density function such that:
@@
    $f\in C^{5},~{}~{}f(0)=1~{}~{},f^{\prime}(0)=0~{}~{}\text{and}~{}~{}\operatorname*{arg\,max}_{x}f(x)=0$ (10)
@@
Proof.
(Theorem 5.1)
For the duration of this proof for notational convenience, denote $\beta_{\text{max}}$ simply by $\beta$ as that is the only temperature considered in this analysis.
$\pi_{\beta}(x)\propto\prod_{i=1}^{d}f^{\beta}(x_{i})$ (18)
with independence sampler proposals generated by
$q_{\beta}(x)=\prod_{i=1}^{d}\phi_{\beta}(x_{i}).$
$\mathbb{P}(Accept)=\min\left\{1,e^{B(d)}\right\}$ (19)
where
The logged acceptance ratio, $B$ is given by
$h(x)=h(0)+\frac{x^{2}}{2}H+\frac{x^{3}}{6}h^{\prime\prime\prime}(0)+\frac{x^{4}}{24}h^{{}^{\prime\prime\prime\prime}}(0)+\frac{x^{5}}{120}h^{v}(\xi_{x}).$ (22)
Thus (21) becomes
$~{}~{}~{}~{}~{}~{}~{}~{}+\frac{\beta}{120}\left(y_{i}^{5}h^{v}(\xi_{y_{i}})-x_{i}^{5}h^{v}(\xi_{x_{i}})\right)\Bigg{]}.$
Lemma 5.2.
is such that as $d\rightarrow\infty$
$W(d)\Rightarrow N\left(-\frac{5h^{\prime\prime\prime}(0)^{2}}{12\ell(-h^{\prime\prime}(0))^{3}},\frac{5h^{\prime\prime\prime}(0)^{2}}{6\ell(-h^{\prime\prime}(0))^{3}}\right)$ (24)
where the convergence is weak convergence.
Proof.
See LABEL:proofRegular. ∎
Lemma 5.3.
Proof.
See LABEL:proofRegular. ∎
Lemma 5.4.
Under the setting of Theorem 5.1, then if $\beta=\ell d$ as in (28) then
$T_{2}(d):=\sum_{i=1}^{d}\frac{\beta}{120}\left(y_{i}^{5}h^{v}(\xi_{y_{i}})-x_{i}^{5}h^{v}(\xi_{x_{i}})\right)$
is such that $T_{2}(d)\rightarrow 0$ in $\mathbb{P}$ as $d\rightarrow\infty$.
Proof.
See LABEL:proofRegular. ∎
Taking the results of Lemmata 5.2, 5.3 and 5.4 and trivial application of Slutsky’s Theorem, then we have that as $d\rightarrow\infty$ then
$B(d)=W(d)+T(d)\Rightarrow B\sim N\left(-\frac{5h^{\prime\prime\prime}(0)^{2}}{12\ell(-h^{\prime\prime}(0))^{3}},\frac{5h^{\prime\prime\prime}(0)^{2}}{6\ell(-h^{\prime\prime}(0))^{3}}\right)$ (25)
where the convergence is weak convergence.
Proposition 1.
Suppose that $G\sim N(-\frac{\sigma^{2}}{2},\sigma^{2})$,
Proof.
A routine calculation (see e.g., Roberts et al. (1997)). ∎
@@ 5.2 Relaxing the regularity conditions
The previous section established the primary result under certain smoothness and regularity conditions on the target density. Yet, it remains an open question whether these assumptions can be loosened. While the $C^{2}$ smoothness at the mode is necessary for weak convergence to a Gaussian distribution, Theorem 5.1 also reveals the critical role of the third derivative. If the third derivative of the log-density were zero at the mode, the Mode Leap Point Independence Sampler could achieve an asymptotically optimal acceptance rate, namely $\lim_{d\to\infty}a(d)=1$. Given that the third derivative quantifies skewness—a key feature that annealing is designed to correct—one may naturally ask what transpires if its continuity at the mode is relaxed.
@@
    $f\in\mathcal{C}^{2}{\left(\mathbb{R}\right)}\cap\mathcal{C}^{5}{\left(\mathbb{R}^{*}\right)}\>,\>f{\left(0\right)}=1\>,\>f^{\prime}{\left(0\right)}=0\>,\>\operatorname*{arg\,max}{f{\left(x\right)}}=0$ (26)
@@
    Assume that
    $\forall k\in\{3,4,5\},\>f^{\left(k\right)}\text{ can be extended by continuity on }\mathbb{R}_{+}\text{ and on }\mathbb{R}_{-}$ (27)
@@
Remark 3.
One can independently observe the correspondence between the extended result in Theorem 5.5 and the initial findings of Theorem 5.1 when the marginal iid components $f$ are chosen such that the third derivative of $h$ is continuous at 00.
@@
See LABEL:proofRelaxed. ∎
Theorem 5.5 establishes that discontinuities in the third derivative of the target distribution at its modes do not prevent the Mode Leap Point Independence Sampler from achieving a positive acceptance rate in the asymptotic regime, when the inverse temperature scales in $\mathcal{O}(d)$. For a more in-depth investigation of ALPS under less restrictive smoothness and regularity assumptions, readers are referred to Queniat (2024, Sections 3 and 4).
@@ 6 Computational Complexity of ALPS
We refer the reader to Roberts et al. (2022) for a precise description of these results. For our purposes here, we note that these results suggest that the full version of ALPS has a mixing time that scales as $\mathcal{O}(d)$ and that if the QuanTA transformation-aided swap moves are not deployed (in favour of standard parallel tempering swap moves) then the mixing time scales as $\mathcal{O}(d[\log(d)]^{2})$.
@@ 7 Applications of ALPS
7.1 Example 1: Synthetic Hard 20-Dimensional Problem
@@
$\pi(x)\propto\sum_{k=1}^{13}\prod_{j=1}^{20}\omega_{k}\;\phi\left(\frac{x_{j}-(\mu_{k})_{j}}{\sigma_{k}}\right)\Phi\left(\alpha\frac{x_{j}-(\mu_{k})_{j}}{\sigma_{k}}\right)$
where the mixing weights $\omega_{k}>0\ \forall k\in\{1,\ldots,13\}$, $\sum_{k}\omega_{k}=1$, skewness parameter $\alpha=10$, and where $\phi(\cdot)$ and $\Phi(\cdot)$ denote the density and CDF of a standard normal distribution, respectively. The mode locations $\vec{\mu}_{k}$ are distributed in a similar manner to the 13-mode example in Tjelmeland and Hegstad (2001), except in a 20-dimensional hypercube instead of only 2 dimensions. This example has been designed to illustrate the problem with torpid mixing of parallel tempering (Woodard et al., 2009a) due to the heterogeneity of the scales $\sigma_{k}$ of the mixture components.
@@
The LAIS only uses a single temperature level which is the target level. For ALPS, 7 temperature levels were used for the annealing part of the algorithm with a cooling schedule given by powers in the vector $(1,4,16,64,256,1024,4096)$. The maximum level of cooling was excessively cold as non-degenerate acceptances for mode jumping moves were achieved at warmer temperatures (in agreement with the results of Section 5). However, with the inter-temperature mixing proving extremely fast at the colder temperatures and the target distribution evaluations being very cheap due to the synthetic nature of the test then one could afford to go much cooler to get high acceptance rates for the mode jumping moves. In this example the mode jumping acceptance rates were around 0.85.
In contrast, PT required 23 temperature levels which were geometrically spaced with common ratio of 0.6. This gave the suggested optimal temperature swap rates of 0.234 between consecutive temperature level pairs along with a temperature that appeared hot enough to explore the entire state space. Note that the minimum power $\beta_{min}=0.6^{22}\approx 1.3\times 10^{-5}$ used for PT is slightly hotter than the inverse temperature used for mode-finding in ALPS, $\beta_{EC}=2.5\times 10^{-5}$, in order to make a fair comparison between the two algorithms. Even so, the PT algorithm fails to find all 13 modes even after 500,000 iterations. This is due to the differences in scale $\sigma_{k}$ between the mixture components, as predicted by Woodard et al. (2009a). Due to torpid mixing, it could take several million more iterations for PT to converge.
The most difficult aspect of tuning ALPS was with respect to the Exploration Component for mode-finding. This was also highlighted in the real-data example in the following section. As mentioned previously, finding “narrow modes” is never guaranteed in the finite run of the algorithm if their location is unknown (Fong et al., 2019). In both LAIS and ALPS where mode-finding is required, it was found that an inverse temperature of $\beta_{EC}=0.000025$ worked very well and rapidly found the modes; typically within a few thousand iterations of the algorithm. However, in this example the number of modes was known and so the mode finding could be optimised for ALPS and LAIS; something not possible in a real data example where one may continue mode searching throughout the duration or even doing some pre-computation dedicated to mode searches (as was the case in the following real-data example in Section 7.2).
@@
In addition to the simulation study presented here, we have also included a performance comparison of ALPS, LAIS and PT across 10 repeated runs of each algorithm in LABEL:sec:Ex3. This example has 4 modes instead of 13 and we run each algorithm for 200,000 iterations, discarding the first 20,000 as burn-in.
@@ 7.2 Example 2: Seemingly-Unrelated Regression Model
where $\vec{y}_{m}$ is a column vector of $N_{m}$ observations for the $m$th response, $\vec{\theta}_{m}$ is a vector of $J_{m}$ regression coefficients, $\mathbf{X}_{m}$ is an $N_{m}\times J_{m}$ design matrix of covariates, and $\vec{\epsilon}_{m}$ is a vector of residual errors. For notational convenience, we assume that $N_{1}=N_{2}=\dots=N_{M}=N$ and likewise that $J_{1}=J_{2}=\dots=J_{M}=J$, although in full generality that is not always the case. For example, Kmenta (1986, p. 687) included a SUR model with $J_{1}=2$ and $J_{2}=3$ covariates.
The profile log-likelihood of the SUR model is thus:
Due to the quadratic form of the SUR model, it was long assumed that its likelihood was log-concave and therefore the iterative estimation procedure of Zellner (1962) was guaranteed to converge to the global maximum. For example, see Srivastava and Giles (1987, p. 157) and Greene (1997, p. 686). However, in fact these models are an example of a curved exponential family (Sundberg, 2010). Drton and Richardson (2004) showed in the bivariate case that maximising (31) is equivalent to finding the real roots of a fifth-degree polynomial. The profile log-likelihood preserves all stationary points of the log-likelihood, so if the likelihood is multi-modal then so will $\ell(\mathbf{Y}\mid\bm{\theta})$ (Drton and Richardson, 2004, Lemma 1). This means that the likelihood could have one, two, or three local maxima. Drton and Richardson (2004) provided a specific example of a dataset where the SUR model exhibits two local modes, which we will examine in more detail in the following section.
@@ 7.2.1 Bimodal Example
Figure 3: (a) 3D perspective plot of the profile likelihood (31) for the bivariate SUR model, adapted from Drton and Richardson (2004, Fig. 1); (b) Annealed likelihood at the coldest temperature level used in ALPS.
In the simplest, bivariate case with $M=2$ responses and only $J=1$ covariates per response, the model (30) can be written as:
@@
Drton and Richardson (2004) featured an example dataset with $N=8$ observations. The profile likelihood (31) is illustrated in Figure 3. There are two modes at $\bm{\theta}=$ (0.78, 1.54) and (2.76, 2.50) and a saddle point at (1.62, 2.03). The iterative algorithm is initialised at the ordinary least squares estimate, $\hat{\bm{\theta}}=$ (1.25, 1.78). Using the R package ‘systemfit’ (Henningsen and Hamann, 2007) with a tolerance level of $10^{-6}$, the algorithm of Zellner (1962) converges to the first mode after 25 iterations. However, the profile log-likelihood at this mode is -27.72, whereas at the other mode it is -27.35. This algorithm has therefore failed to find the global maximum.
@@
We run ALPS using the same profile log-likelihood function, with three inverse-temperature levels $(1,\sqrt{10},10)$ as well as the hot state 0.5. The Exploration Component locates both modes after 236 iterations, as shown in Figure 4 (d), and ALPS takes 2.6 seconds for 10,000 iterations in total. The MCMC trace plots for the two parameters at the coldest temperature are shown in Figures 4 (a) and 4 (b), respectively, and a histogram for $\theta_{1}$ is shown in Figure 4 (c). The jump rate for mode-hopping proposals was 0.9 at the coldest temperature, with acceptance rates of 0.067 for the EC, 0.265 for within-temperature moves and 0.331 for position-dependent moves at the target temperature. The temperature-swapping acceptance rate was 0.508 between the first two temperatures, and 0.847 at the coldest temperature.

### v3

@@ 5.1 Optimal Scaling of the Coldest Temperature Level
Roberts et al. (2022) demonstrates that, at least under stylised conditions, if we can ensure that jumping between modes is possible at the coldest temperature, then the entire algorithm (suitably time-scaled) converges to a process (Walsh Brownian Motion) which itself has a mixing time which is independent of $d$. (While the exact result (Theorem 4) in that paper assumes that jumping between modes at the cold temperature always occurs, it is easy to see from the proof that it extends routinely to just requiring that jumps happen with probability which are bounded away from 0.) Therefore this section will be dedicated to demonstrating that at a sufficiently cold inverse temperature (we need $\beta=\mathcal{O}(d)$) then the probability of accepting a move to a different mode in the cold temperature is bounded away from 0.
@@ 5.1.1 Assumptions on the Target Density
where $\sigma_{ij}\in\mathbb{R}_{+}$, $\mu_{ij}\in\mathbb{R}$ and $f$ is an un-normalised univariate density function such that:
@@
    $f\in\mathcal{C}^{5},~~f(0)=1~~,f^{\prime}(0)=0~~\text{and}~~\operatorname*{arg\,max}_{x}f(x)=0$ (10)
@@
Notice that the tempered distribution in (9) is well approximated by the HAT distributions when the modes become well separated, a regime that naturally arises as the inverse temperature $\beta$ increases. These assumptions are therefore natural and well suited to the asymptotic regime of interest.
@@
The limiting acceptance probability in Theorem 5.1 depends explicitly on the local asymmetry through the third derivative of the log-density at the mode. Specifically, the quantity
$\gamma\;:=\;-\,\frac{h^{\prime\prime\prime}(0)}{\bigl(-h^{\prime\prime}(0)\bigr)^{3/2}}$ (18)
represents the local skewness of the target density around its mode. This standardized third derivative arises naturally in third-order Laplace and saddlepoint approximations and quantifies departures from local Gaussian behaviour; see, for example, Daniels (1954). From a Bayesian perspective, such skewness occurs naturally in posteriors arising from asymmetric or non-Gaussian priors, for instance a Gamma prior combined with a Poisson likelihood. More generally, local skewness reflects an intrinsic asymmetry of the distribution in a neighbourhood of its mode, whereby probability mass accumulates unevenly on either side of the mode.
Theorem 5.1 shows that, even as the magnitude $|\gamma|$ increases and the local asymmetry becomes more pronounced, the Mode Leap Point Independence sampler retains a strictly positive asymptotic acceptance probability under linear scaling of the coldest temperature level with dimension. Moreover, any target acceptance probability in $(0,1)$ can be achieved, under linear scaling $\beta_{\max}=\ell d$, by selecting a sufficiently large constant $\ell$, with higher skewness requiring a larger value of this constant while preserving the same linear scaling with dimension. This shows that ALPS can accommodate strong skewness without altering its mixing capabilities.
@@ 5.2 Relaxing the regularity conditions
The previous section established the primary result under certain smoothness and regularity conditions on the target density. Yet, it remains an open question whether these assumptions can be loosened. While $C^{2}$ smoothness at the mode is necessary to ensure weak convergence to a Gaussian distribution, Theorem 5.1 highlights the central role played by the local skewness coefficient (18) at the mode, in determining the performance of the algorithm. In the favourable case where this coefficient vanishes, the local behaviour of the target is effectively symmetric, and the $\beta_{\max}=\mathcal{O}(d)$ scaling becomes more than sufficient, yielding an asymptotic acceptance probability satisfying $\lim_{d\to\infty}a(d)=1$. This suggests that in the absence of local skewness, a substantially milder annealing regime—potentially sublinear in dd—may already suffice to maintain a non-vanishing acceptance rate for the Mode Leap Point Independence sampler. This naturally raises the question of what happens when skewness arises not from a smooth third derivative, but instead from piecewise-smooth behaviour, leading to a discontinuity of the third derivative at the mode.
@@
    $f\in C^{2}(\mathbb{R}),\quad f$ $\in C^{5}((-\infty,0))\cap C^{5}((0,\infty)),$ (19)
    $f(0)=1,\quad f^{\prime}(0)$ $=0,\quad\operatorname*{arg\,max}f(x)=0,$
@@
    Furthermore, assume that [^1] We assume continuity at 0 up to the second derivative. Higher-order derivatives are assumed continuous only on $\mathbb{R}_{\pm}$; the corresponding one-sided limits at 0 are allowed to differ.
[…]
@@
See Appendix A.2. ∎
Theorem 5.2 establishes that discontinuities in the third derivative of the target distribution at its modes do not prevent the Mode Leap Point Independence Sampler from achieving a positive acceptance rate in the asymptotic regime, provided $\beta_{\text{max}}$ scales as $\mathcal{O}(d)$. In this piecewise-smooth setting, the asymptotic acceptance probability decomposes into a symmetric contribution, corresponding to the overall magnitude of local skewness on either side of the mode, and an asymmetric contribution induced by the jump in the third derivative, which captures the mismatch between the left and right local geometries. In particular, when the third derivative is continuous at the mode, the skewness jump vanishes and the result reduces to the smooth-case acceptance of Theorem 5.1.
Such piecewise-smooth behaviour arises in Bayesian settings where the log-posterior is constructed from components that are smooth on each side of the mode but differ in higher-order behaviour, for example due to asymmetric or piecewise-defined priors, truncated or constrained parameter spaces. For a more in-depth investigation of ALPS under less restrictive smoothness and regularity assumptions, readers are referred to Queniat (2024, Sections 3 and 4).
@@ 6 Computational Complexity of ALPS
We refer the reader to Section 5 and in particular Corollaries 2 and 3 of Roberts et al. (2022) for a detailed description of these results (though note the typo in Corollary 3 which should be clear on reading the previous paragraph). For our purposes here, we note that these results suggest that the full version of ALPS has a mixing time that scales as $\mathcal{O}(d)$ and that if the QuanTA transformation-aided swap moves are not deployed (in favour of standard parallel tempering swap moves) then the mixing time scales as $\mathcal{O}(d[\log(d)]^{2})$.
@@ 7 Applications of ALPS
7.1 A Synthetic 13-Modal Benchmark Distribution
7.1.1 Example 1: The 20-Dimensional Case
@@
$\pi(x)\propto\sum_{k=1}^{13}\omega_{k}\prod_{j=1}^{20}\phi\left(\frac{x_{j}-(\mu_{k})_{j}}{\sigma_{k}}\right)\Phi\left(\alpha\frac{x_{j}-(\mu_{k})_{j}}{\sigma_{k}}\right)$
where the mixing weights $\omega_{k}>0\ \forall k\in\{1,\ldots,13\}$, skewness parameter $\alpha=10$, and where $\phi(\cdot)$ and $\Phi(\cdot)$ denote the density and CDF of a standard normal distribution, respectively. The mode locations $\vec{\mu}_{k}$ are distributed in a similar manner to the 13-mode example in Tjelmeland and Hegstad (2001), except in a 20-dimensional hypercube instead of only 2 dimensions. This example has been designed to illustrate the problem with torpid mixing of parallel tempering (Woodard et al., 2009a) due to the heterogeneity of the scales $\sigma_{k}$ of the mixture components.
@@
The LAIS only uses a single temperature level which is the target level. For ALPS, 7 temperature levels were used for the annealing part of the algorithm with a coldest inverse temperature $\beta_{\text{max}}=4096$ and a cooling schedule given by powers in the vector $(1,4,16,64,256,1024,4096)$. The maximum level of cooling was excessively cold as non-degenerate acceptances for mode jumping moves were achieved at warmer temperatures (in agreement with the results of Section 5). However, because mixing across temperatures proves extremely quick at colder temperatures and evaluating the synthetic target is inexpensive, it is possible to use much lower temperatures to get higher acceptance rates for the mode-jumping moves. In this example the mode jumping acceptance rates were around 0.85.
In contrast, PT required 23 temperature levels which were geometrically spaced with common ratio of 0.6. This gave the suggested optimal temperature swap rates of 0.234 between consecutive temperature level pairs along with a temperature that appeared hot enough to explore the entire state space. Note that the minimum power $\beta_{min}=0.6^{22}\approx 1.3\times 10^{-5}$ used for PT is slightly hotter than the inverse temperature used for mode-finding in ALPS, $\beta_{\text{hot}}=2.5\times 10^{-5}$, in order to make a fair comparison between the two algorithms. Even so, the PT algorithm fails to find all 13 modes even after 500,000 iterations. This is due to the differences in scale $\sigma_{k}$ between the mixture components, as predicted by Woodard et al. (2009b). Due to torpid mixing, it could take several million more iterations for PT to converge.
In all three algorithms random walk Metropolis was utilised for the within temperature moves; in the case of ALPS and LAIS the random walk Metropolis moves used pre-conditioned covariance structure informed by the covariance matrix estimated at the local mode point by the Laplace approximation. In all cases the within temperature moves were tuned to have the suggested optimal acceptance rate of 0.234.
The most difficult aspect of tuning ALPS was with respect to the Exploration Component for mode-finding. This was also highlighted in the real-data example in the following section. As mentioned previously, finding “narrow modes” is never guaranteed in the finite run of the algorithm if their location is unknown (Fong et al., 2019). In both LAIS and ALPS where mode-finding is required, it was found that an inverse temperature of $\beta_{\text{hot}}=0.000025$ worked very well and rapidly found the modes; typically within a few thousand iterations of the algorithm. However, in this example the number of modes was known and so the mode finding could be optimised for ALPS and LAIS; something not possible in a real data example where one may continue mode searching throughout the duration or even doing some pre-computation dedicated to mode searches (as was the case in the following real-data example in Section 7.2).
@@
In addition to the simulation study presented here, we have also included a performance comparison of ALPS, LAIS and PT across 10 repeated runs of each algorithm in Appendix D. This example has 4 modes instead of 13 and we run each algorithm for 200,000 iterations, discarding the first 20,000 as burn-in.
7.1.2 Scaling with Dimension at the Cold Level
We now empirically validate the dimensional scaling result of Theorem 5.1 for the coldest temperature level using the synthetic multimodal example introduced above. The aim of this experiment is to assess whether the asymptotic acceptance rate derived in Theorem 5.1 is verified in high-dimensional settings.
Given a prescribed target acceptance rate $a\in(0,1)$ at the cold level, Theorem 5.1 implies that the linear scaling
$\beta_{\text{max}}=\ell(a)\,d$ (22)
yields a non-degenerate asymptotic acceptance rate. Inverting the limiting expression in Theorem 5.1, the constant $\ell(a)$ in (22) is given explicitly by
$\ell(a)=\frac{5h^{\prime\prime\prime}(0)^{2}}{24\,(-h^{\prime\prime}(0))^{3}\,\left[\Phi^{-1}(a/2)\right]^{2}},$ (23)
where $h(x)=\log f(x)$ and $\Phi$ denotes the standard Gaussian distribution function. In this synthetic example, the function $f$ is fully specified, so that (23) can be computed explicitly.
Using this expression, we compute $\beta_{\text{max}}$ for increasing dimensions and apply the Mode Leaping Independence Sampler to the same synthetic mixture model, keeping the number of modes fixed. In this experiment, the skewness parameter is set to $\alpha=5.0$. For each dimension, we report the empirical acceptance rate of the mode-jumping proposals.
The results are shown in Figure 3. As the dimension increases, the empirical acceptance rates approach the prescribed target value, providing empirical support for the linear scaling $\beta_{\text{max}}=\ell(a)\,d$, thereby verifying the asymptotic result stated in Theorem 5.1.
Figure 3: Empirical acceptance rates of the mode-jumping proposals at the coldest temperature level as a function of dimension, averaged over $128$ runs of $16384$ iterations. For each dimension $d$, the cold inverse temperature is set to $\beta_{\text{max}}=\ell(a)\,d$, where $\ell(a)$ is computed according to (23). Target acceptance rates $a$ are set to (a) 0.3, (b) 0.5, (c) 0.7 and (d) 0.9.
@@ 7.2 Example 2: Seemingly-Unrelated Regression Model
where $\vec{y}_{m}$ is a column vector of $N_{m}$ observations for the $m$th response, $\vec{\theta}_{m}$ is a vector of $J_{m}$ regression coefficients, $\mathbf{X}_{m}$ is an $N_{m}\times J_{m}$ design matrix of covariates, and $\vec{\epsilon}_{m}$ is a vector of residual errors. For notational convenience, we assume that $N_{1}=N_{2}=\dots=N_{M}=N$ and likewise that $J_{1}=J_{2}=\dots=J_{M}=J$, although in full generality that is not always the case. See Appendix E for further introductory information on the SUR model.
Drton and Richardson (2004) provided a specific example of a dataset where the SUR model exhibits two local modes, which we will examine in more detail in the following section.
@@ 7.2.1 Bimodal Example
In the simplest, bivariate case with $M=2$ responses and only $J1$ covariates per response, the model can be written as:
@@
Drton and Richardson (2004) featured an example dataset with $N=8$ observations. The profile likelihood, derived in Appendix E, is illustrated in Figure 4. There are two modes at $\bm{\theta}=$ (0.78, 1.54) and (2.76, 2.50) and a saddle point at (1.62, 2.03). The iterative algorithm is initialised at the ordinary least squares estimate, $\hat{\bm{\theta}}$= (1.25, 1.78). Using the R package ‘systemfit’ (Henningsen and Hamann, 2007) with a tolerance level of $10^{-6}$, the algorithm of Zellner (1962) converges to the first mode after 25 iterations. However, the profile log-likelihood at this mode is -27.72, whereas at the other mode it is -27.35. This algorithm has therefore failed to find the global maximum.
Figure 4: (a) 3D perspective plot of the profile likelihood for the bivariate SUR model, adapted from Drton and Richardson (2004, Fig. 1); (b) Annealed likelihood at the coldest temperature level used in ALPS.
@@
We run ALPS using the same profile log-likelihood function, with $\beta_{\text{max}}=10$ and a geometric cooling schedule composed of three inverse-temperature levels $(1,\sqrt{10},10)$ as well as the hot state 0.5. The Exploration Component locates both modes after 236 iterations, as shown in Figure 5 (d), and ALPS takes 2.6 seconds for 10,000 iterations in total. The MCMC trace plots for the two parameters at the coldest temperature are shown in Figures 5 (a) and 5 (b), respectively, and a histogram for $\theta_{1}$ is shown in Figure 5 (c). The jump rate for mode-hopping proposals was 0.9 at the coldest temperature, with acceptance rates of 0.067 for the EC, 0.265 for at the intermediate temperature and 0.331 at the target temperature. The temperature-swapping acceptance rate was 0.508 between the first two temperatures, and 0.847 at the coldest temperature.